# A resource-efficient variational quantum algorithm for mRNA codon optimization

### v1

@@ 5 Conclusion
This research is substantial contribution to the exploration of quantum computing for mRNA codon optimization. The thoughtful recognition of potential areas for refinement underscores the depth of the research. Our resource-efficient quantum algorithm halves the required number of qubits in comparison to the existing approach [3].
@@
  • Polypeptide are a group of polymers made from a chain of amino acids.
@@
We leverage a dense encoding method to represent each synonymous codon for each amino acid in a protein chain. This method uses a string of qubits with each qubit assigned a value of ‘1’ or ‘0’ to define mRNA sequence, while each synonymous codon has its own binary qubit to be represented. The number of required qubits for each amino acid is related to the number of synonymous codon available (equation 2). Equation 5 represents the complete mRNA sequence for a protein chain, where $i$ represents the $i$th amino acid in a protein, and k represents the kth qubits. The choice of encoding method can significantly impact the optimization process and the resulting mRNA sequences.
@@
The critical part of mRNA codon optimization involves the construction of the Hamiltonian, a fundamental element in the optimization process. The Hamiltonian takes into consideration several key factors that are essential for optimizing codon in a given protein sequence. These factors encompass codon usage bias, target GC concentration, codon context, and the presence of sequentially repeated nucleotides [3]. Additionally, the Hamiltonian accounts for constraint items related to redundant encoding, as they cannot be translated to an actual mRNA sequence.

### v2

@@ 5 Conclusion
This research is a substantial contribution to the exploration of quantum computing for mRNA codon optimization. The thoughtful recognition of potential areas for refinement underscores the depth of the research. Our resource-efficient quantum algorithm halves the required number of qubits in comparison to the existing approach [3].
@@
  • Polypeptides are a group of polymers made from a chain of amino acids.
@@
We leverage a dense encoding method to represent each synonymous codon for each amino acid in a protein chain. This method uses a string of qubits with each qubit assigned a value of ‘1’ or ‘0’ to define the mRNA sequence, while each synonymous codon has its own binary qubit to be represented. The number of required qubits for each amino acid is related to the number of synonymous codons available (equation 2). Equation 4 represents the complete mRNA sequence for a protein chain, where $i$ represents the $i$th amino acid in a protein, and k represents the kth qubits. The choice of encoding method can significantly impact the optimization process and the resulting mRNA sequences.
@@
The critical part of mRNA codon optimization involves the construction of the Hamiltonian, a fundamental element in the optimization process. The Hamiltonian takes into consideration several key factors that are essential for optimizing codons in a given protein sequence. These factors encompass codon usage bias, target GC concentration, codon context, and the presence of sequentially repeated nucleotides [3]. Additionally, the Hamiltonian accounts for constraint items related to redundant encoding, as they cannot be translated to an actual mRNA sequence.